Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed leave it blank. If no reaction occurs leave all boxes blank and click on "submit" Predict if a reaction will occur when Cu(s) and hydroiodic acid(aq) are combined. If a reaction occurs, write a balanced equation for the reaction. Retry Entire Group 3 more group attempts remaining Submit Answer
Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed leave it blank If no reaction occurs leave all boxes blank and click on "submit". Use the table on the information page to predict if a reaction will occur when Mn(s) and hydrobromic acid(aq) are combined. If a reaction occurs, write a balanced equation for the reaction. Retry Entire Group 3 more group attempts remaining Submit Answer
Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed leave it blank If no reaction occurs leave all boxes blank and click on "Submit" Write a net ionic equation for the reaction that occurs when aqueous solutions of barium hydroxide and hydrochloric acid are combined. (Use H' instead of Hy0t) Submit Answer Retry Entire Group 1 more group attempt remaining
Use the pull-down boxes to specify states such as (aq) or (s). If a box is not needed leave it blank. If no reaction occurs leave all boxes blank and click on "Submit" Write a net ionic equation for the reaction that occurs when aqueous solutions of sodium hydroxide and hypochlorous acid are combined 1 more group attempt remaining Retry Entire Group Submit Answer
1. How many GRAMS of cobalt(II) acetate are present in 1.90 moles of this compound ? grams. 2. How many MOLES of cobalt(I) acetate are present in 2.98 grams of this compound? moles Submit Answer Retry Entire Group 2 more group attempts remaining
Use the References to access important values if needed for this question 1. How many MOLES of phosphorus trichloride are present in 3.94 grams of this compound ? moles 2. How many GRAMS of phosphorus trichloride are present in 4.13 moles of this compound ? grams Submit Answer Retry Entire Group 2 more group attempts remaining
Answer #1

1. The reaction between Cu and Hydroiodic acid will not occur.

Beacuse Iodine is above the copper in a metal reactivity series. Ie the standard reduction potential of Cu (+0.337) is greater than Hydrogen(0)

Cu2+ + 2e-→ Cu E0red = + 0.337 V

2H^{+} + 2e^{-} \rightarrow H_{2} E0red = 0 V

so copper cannot replace the hydrogen in HI to form CuI2.

2. The reaction between Mn and Hydrobromic acid will occur. Beacause Mn is above the Bromine in metal reactivity series.

Mn_{(s)} + 2HBr _{(aq)} \rightarrow MnBr_{2}_{(aq)} + H_{2}_{(g)}  

3.

Balanced equation  Ba(OH)2(ag)+2HClag) Bal2(ag)2H20

Ionic equation Ba^{2+}_{(aq)} + 2(OH)^{-}_{(aq)} + 2H^{+}_{(aq)} + 2Cl^{-}_{(aq)} \rightarrow Ba^{2+} + 2Cl^{-}_{(aq)} + 2 H_{2}O_{(l)}   

Net ionic equation  2(OH) (aq) + 2H (aq)-ל2H20 (1)

4. Balance equation NaOH_{(aq)} + HClO_{(aq)} \rightarrow NaOCl_{(aq)} + H_{2}O_{(l)}

Ionic equation  dq

net ionic equation  (OH)^{-}_{(aq)} + H^{+}_{(aq)} \rightarrow H_{2}O_{(l)}

Formulae: -

noss Moles = molarmaSS

Mass = Moles \times molar mass

5.

Molar mass of Cobalt(II) acetate = 177.02 g/mol

​​​​​a) mole of cobalt (II) acetate = 1.90 mol

therefore,

Mass = Moles \times molar mass

Mass = 1.911101 × 177.029/mol = 336.34g

b) mass of cobalt(II) acetate = 2.98 g

Therfore,

noss Moles = molarmaSS

Moles = \frac{2.98 g}{177.02 g/mol} = 0.0168 mol

6. Molar mass of phosphorous trichloride = 137.33 g/mol

a) mass of phosphorous trichloride = 3.94 g

Therefore,

noss Moles = molarmaSS

3.94g 0.0287mol Moles - 137.33g mol

b) moles of phosphorous trichloride = 4.13 mol

Therefore,

Mass = Moles \times Molar mass

Mass-4.13mol × 137.33g/mol 567.173g